OpenStudy (anonymous):

Calculate the scalar projection and projection of (5,3) onto (7,-2)

OpenStudy (amistre64):

 cant recall the def of a "scalar projection" to clearly tho

OpenStudy (amistre64):

essentially you take the unit of v (the onto vector) and multiply it by the |u| cos(a), where a is the angle between them

OpenStudy (anonymous):

Ok, so how would you calculate it through the coordinates they gave me?

OpenStudy (amistre64):

\[cos(\alpha)=\frac{u\cdot v}{|u|~|v|}\] \[|u|~cos(\alpha)=\cancel{|u|}\frac{u\cdot v}{\cancel{|u|}~|v|}=|proj_uV|\]

OpenStudy (amistre64):

u = (5,3) ; length of u v = (7,-2) ; length of v --------- dot them we will need to find these properties in order to plug them into the formula

OpenStudy (amistre64):

hmmm, from the google i get that |u| cos(a) is defined as the SCALAR projection, the magnitude of proj_u V

OpenStudy (anonymous):